OpenStudy (anonymous):

Represent the complex number graphically and find the standard form of the number. 8(cos225 + isin225)

OpenStudy (amoodarya):

 \[8(\frac{ -\sqrt{2} }{ 2 }+i\frac{ -\sqrt{2} }{ 2 } )\]

OpenStudy (anonymous):

Thanks! I appreciate your help!